Solve the system of equations and express the answer as an ordered
pair of the form (x,y) or write NO SOLUTION if applicable:
x + 2y = 5
2x - 2y = 4

Answer by palanisamy(496)

The given equatins are
x + 2y = 5 ...(1)
2x - 2y = 4 ...(2)
(1)+(2)=> 3x = 9
x = 9/3
x = 3
Substituting in (1), we get
3+2y = 5
2y = 5-3
2y = 2
y = 2/2
y = 1
Therefore the solution is(3,1)